# TopoTune: A Framework for Generalized Combinatorial Complex Neural Networks

### v3

@@ Abstract
Graph Neural Networks (GNNs) excel in learning from relational datasets, processing node and edge features in a way that preserves the symmetries of the graph domain. However, many complex systems—such as biological or social networks—involve multiway complex interactions that are more naturally represented by higher-order topological domains. The emerging field of Topological Deep Learning (TDL) aims to accommodate and leverage these higher-order structures. Combinatorial Complex Neural Networks (CCNNs), fairly general TDL models, have been shown to be more expressive and better performing than GNNs. However, differently from the GNN ecosystem, TDL lacks a principled and standardized framework for easily defining new architectures, restricting its accessibility and applicability. To address this issue, we introduce Generalized CCNNs (GCCNs), a novel simple yet powerful family of TDL models that can be used to systematically transform any (graph) neural network into its TDL counterpart. We prove that GCCNs generalize and subsume CCNNs, while extensive experiments on a diverse class of GCCNs show that these architectures consistently match or outperform CCNNs, often with less model complexity. In an effort to accelerate and democratize TDL, we introduce TopoTune, a lightweight software for defining, building, and training GCCNs with unprecedented flexibility and ease.
@@ 1 Introduction
To date, research in TDL has largely progressed by taking existing GNNs architectures (convolutional, attentional, message-passing, etc.) and generalizing them one-by-one to a specific TDL counterpart, whether that be on hypergraphs (Feng et al., 2019; Chen et al., 2020a; Yadati, 2020), on simplicial complexes (Roddenberry et al., 2021; Yang & Isufi, 2023; Ebli et al., 2020; Giusti et al., 2022a; Battiloro et al., 2024; Bodnar et al., 2021b; Maggs et al., 2024; Lecha et al., 2025), on cell complexes (Hajij et al., 2020; Giusti et al., 2022b; Bodnar et al., 2021a), or on combinatorial complexes (Battiloro et al., 2025; Eitan et al., 2024). Although overall valuable and insightful, such a fragmented research trend is slowing the development of standardized methodologies and software for TDL, as well as limiting the analysis of its cost-benefits trade-offs (Papamarkou et al., 2024). We argue that these two relevant aspects are considerably hindering the use and application of TDL beyond the community of experts.
@@
TopoX (Hajij et al., 2024b) and TopoBenchmark (Telyatnikov et al., 2024) have become the reference Python libraries for developing and benchmarking TDL models, respectively. However, despite their potential in defining and implementing novel standardized methodologies in the field, the current focus of these packages is on replicating and analyzing existing message-passing CCNNs. Works like Jogl et al. (2022b, a) have instead focused on making TDL accessible and reproducible by porting models to the graph domain. They do so via principled transformations from combinatorial topological domains to graphs. However, although these architectures over the resulting graph-expanded representations are as expressive as their TDL counterparts (using the Weisfeiler-Lehman criterion (Xu et al., 2019b)), they are neither formally equivalent to nor a generalization of their TDL counterparts. Due to collapse of rank information during the graph expansion, the GNNs on the resulting graph do not preserve the same topological information as their TDL counterparts.
@@
Contributions. This works seeks to accelerate TDL research and increase its accessibility and standardization for outside practitioners. To that end, we introduce a novel joint methodological and software framework that easily enables the development of new TDL architectures in a principled way—overcoming the limitations of existing works. We outline our main contributions and specify which of the field’s open problems (as defined in Papamarkou et al. (2024)) they help answer:
@@
  • Systematic Generalization. We propose the first method to systematically generalize any neural network to its topological counterpart with minimal adaptation. Specifically, we define a novel expansion mechanism that transforms a combinatorial complex into a collection of graphs, enabling the training of TDL models as an ensemble of synchronized models. To our knowledge, this is the first method which is designed to work across many topological domains. (Open problems 6, 11: need for foundational, cross-domain TDL.)
@@
  • General Architectures. Our method induces a novel wide class of TDL architectures, Generalized Combinatorial Complex Networks (GCCNs), portrayed in Fig. 1. GCCNs (i) formally generalize CCNNs, (ii) are cell permutation equivariant, and (iii) are as expressive as CCNNs. (Open problem 9: consolidating TDL advantages in a unified theory.)
@@
  • Implementation. We provide TopoTune, a lightweight PyTorch module for developing GCCNs fully integrated into TopoBenchmark (Telyatnikov et al., 2024). Using TopoTune, practitioners can, for the first time, easily define and iterate upon TDL architectures. (Open problems 1, 4: need for accessible TDL and software.)
@@
  • Benchmarking. Using TopoTune, we create a broad class of GCCNs using four base GNNs and one base Transformer over two combinatorial topological spaces (simplicial and cell complexes). A wide range of experiments on graph-level and node-level benchmark datasets shows GCCNs generally outperform existing CCNNs, often with smaller model sizes. Some of these results are obtained with GCCNs that cannot be reduced to standard CCNNs, further underlining our methodological contribution. We will provide all code and scripts if accepted. (Open problem 3: need for standardized benchmarking.)
@@ 4 Generalized Combinatorial Complex Neural Networks
Proofs are provided in Appendix B.1, B.2, and B.3, respectively.
Given Proposition 4.1, GCCNs allow us to define general TDL models using any neighborhood message function $\omega_{\mathcal{N}}$, such as any GNN. Not only does this framework avoid having to approximate CCNN computations, as is the case in previous works  [^4] These models employ GNNs running on one augmented Hasse graph, i.e. a GCCN that, given a collection of neighborhoods $\mathcal{N_{C}}$, uses a single neighborhood $\mathcal{N}_{tot}$ defined, for a cell $\sigma$, as $\mathcal{N}_{tot}(\sigma)=\bigcup_{\mathcal{N}\in\mathcal{N_{C}}}\mathcal{N}(\sigma)$. (Jogl et al., 2022b, a, 2023), but it also enjoys the same permutation equivariance as regular CCNNs (Proposition 4.2). We show in Appendix C that the resulting time complexity of a GCCN is a compromise between a typical GNN and a CCNN. Differently from the work in (Hajij et al., 2023), the fact that GCCNs can have arbitrary neighborhood message functions implies that non message-passing TDL models can be readily defined. To the best of our knowledge, GCCNs are the only objects in the literature that encompass all the above properties.
@@ 5 TopoTune
Our proposed methodology, together with its resulting GCCNs architectures, addresses the challenge of systematically generating principled, general TDL models. Here, we introduce TopoTune, a software module for defining and benchmarking GCCN architectures on the fly—a vehicle for accelerating TDL research. All code and experiment scripts are provided at github.com/geometric-intelligence/TopoBenchmark.
@@
TopoTune is fully integrated into TopoBenchmark (Telyatnikov et al., 2024), a comprehensive package offering a wide range of standardized methods and tools for TDL. Practitioners can access ready-to-use models, training pipelines, tasks, and evaluation metrics, including leading open-source models from TopoX (Hajij et al., 2024b). In addition, TopoBenchmark features the largest collection of topological liftings currently available—transformations that map graph datasets into higher-order topological domains. Together, TopoBenchmark and TopoTune organize the vast design space of TDL into an accessible framework, providing unparalleled versatility and standardization for practitionners.
@@ 6 Experiments
Table 1: Cross-domain, cross-task, cross-expansion, and cross-$\omega_{\mathcal{N}}$ comparison of GCCN architectures with top-performing CCNNs benchmarked on TopoBenchmark (Telyatnikov et al., 2024). Best result is in bold and results within 1 standard deviation are highlighted blue. Experiments are run with 5 seeds. We report accuracy for classification tasks and MAE for regression.
@@
CCNN (Best Model on TopoBenchmark) 80.43 ± 1.78 76.13 ± 2.70 76.67 ± 1.48 75.35 ± 1.50 0.34 ± 0.01 87.44 ± 1.28 75.63 ± 1.58 88.64 ± 0.36
@@
CCNN (Best Model on TopoBenchmark) 76.17 ± 6.63 75.27 ± 2.14 76.60 ± 1.75 77.12 ± 1.07 0.36 ± 0.02 82.27 ± 1.34 71.24 ± 1.68 88.72 ± 0.50
@@
CCNN (Best Model on TopoBenchmark) 80.43 ± 4.09 76.63 ± 1.74 75.18 ± 1.24 74.93 ± 2.50 0.51 ± 0.01 88.92 ± 0.44 74.93 ± 1.39 89.62 ± 0.25
We present experiments showcasing a broad class of GCCN’s constructed with TopoTune. These models consistently match, outperform, or finetune existing CCNNs, often with smaller model sizes. TopoTune’s integration into the TopoBenchmark experiment infrastructure ensures a fair comparison with CCNNs from the literature, as data processing, domain lifting, and training are homogeonized.
@@ 6.1 Experimental Setup
We generate our class of GCCNs by considering ten possible choices of neighborhood structure $\mathcal{N}_{\mathcal{C}}$ (including both regular and per-rank, see Appendix E.1) and five possible choices of $\omega_{\mathcal{N}}$: GCN (Kipf & Welling, 2017), GAT (Velickovic et al., 2017), GIN (Xu et al., 2019a), GraphSAGE (Hamilton et al., 2017), and Transformer (Vaswani et al., 2017). We import these models directly from PyTorch Geometric (Fey & Lenssen, 2019) and PyTorch (Paszke et al., 2019). TopoTune enables running GCCNs on both an ensemble of strictly augmented Hasse graphs (eq. 6) and an augmented Hasse graph (eq. 5). While CCNN results reflect extensive hyperparameter tuning by Telyatnikov et al. (2024), we largely fix GCCN training hyperparameters using the TopoBenchmark default configuration.
@@
We include a wide range of benchmark tasks (see Appendix E.2). MUTAG, PROTEINS, NCI01, and NCI09 (Morris et al., 2020) are graph-level classification tasks about molecules or proteins. ZINC (Irwin et al., 2012) (subset) is a graph-level regression task about solubility. At the node level, the Cora, CiteSeer, and PubMed tasks (Yang et al., 2016) involve classifying publications within citation networks. We consider two topological domains: simplicial and cellular complexes. We use TopoBenchmark’s lifting processes to infer higher-order relationships in these data. We use node features to construct edge and face features.
@@ 6.2 Results and Discussion
GCCNs outperform CCNNs.
Table 1 portrays a cross-comparison between top-performing CCNN models and our class of GCCNs. GCCNs outperform CCNNs in the simplicial and cellular domains in all datasets. Notably, GCCNs in these domains achieve comparable results to hypergraph CCNNs, a feat unattainable by CCNNs in node-level tasks. Out of the 16 domain/dataset combinations considered in our experiments, GCCNs outperform the best counterpart CCNN by $>1\sigma$ in 11 cases. Representing complexes as ensembles of augmented Hasse graphs, rather than a single augmented Hasse graph, consistently improves results.
Generalizing existing CCNNs to GCCNs improves performance.
TopoTune makes it easy to iterate upon and improve preexisting CCNNs by replicating their architecture in a GCCN setting. For example, TopoTune can generate a counterpart GCCN by replicating a CCNN’s neighborhood structure, aggregation, and training scheme. We show in Table 3 in Appendix F that counterpart GCCNs can achieve comparable or better results than SCCN (Yang et al., 2022) and CWN (Bodnar et al., 2021a) just by sweeping over additional choices of $\omega_{\mathcal{N}}$. In the single augmented Hasse graph regime, GCCN models are consistently more lightweight, up to half their size (see Table 5).
GCCNs perform competitively to CCNNs with fewer parameters.
GCCNs are often more parameter efficient than existing CCNNs in simplicial and cellular domains, and in some instances (MUTAG, NCI1, NCI09), even in the hypergraph domain. We refer to Table 4. Even as GCCNs become more parameter-intensive for large graphs with high-dimensional embeddings—as seen in node-level tasks—they remain competitive. For instance, on the Citeseer dataset, a GCCN ($\omega_{\mathcal{N}}$ = GraphSAGE) outperforms the best existing CCNN while being 28% smaller. Training times in Appendix H show GCCNs train comparably on smaller datasets but slow down on larger ones, likely due to TopoTune’s on-the-fly graph expansion. Preprocessing this expansion in future work could mitigate the lag.
TopoTune finds parameter-efficient GCCNs.
By easily exploring a wide landscape of possible GCCNs for a given task, TopoTune helps identify models that maximize performance while minimizing model size. Fig. 5 illustrates this trade-off by comparing the performance and size of selected GCCNs (see Appendix I for more). On the PROTEINS dataset, two GCCNs using per-rank neighborhood structures (orange and purple) achieve performance within 2% of the best result while requiring as little as 48% of the parameters. This reduction is due to fewer neighborhoods $\mathcal{N}$, resulting in fewer $\omega_{\mathcal{N}}$ blocks per GCCN layer. Similarly, on ZINC, lightweight neighborhood structures (orange and dark green) are competitive with reduced parameter costs. Node-level tasks, see less benefit, likely due to the larger graph sizes and higher-dimensional input features.
@@
Fig. 5 also provides insights into the impact of neighborhood message functions. On ZINC, GIN clearly outperforms all other models, which do not even appear in the plot’s range. In the less clear-cut cases of PROTEINS and Citeseer, we observe a trade-off between neighborhood structure and message function complexity. We find that larger base models (GIN, GraphSAGE) on lightweight neighborhood structures perform comparably to simpler base models (GAT, GCN) on larger neighborhood structures. This tradeoff warrants further research on the dataset-specific importance of neighborhood choice, or lack thereof.

### v4

@@ Abstract
Graph Neural Networks (GNNs) excel in learning from relational datasets as they preserve the symmetries of the graph domain. However, many complex systems—such as biological or social networks—involve multiway complex interactions that are more naturally represented by higher-order topological domains. The emerging field of Topological Deep Learning (TDL) aims to accommodate and leverage these higher-order structures. Combinatorial Complex Neural Networks (CCNNs), fairly general TDL models, have been shown to be more expressive and better performing than GNNs. However, differently from the GNN ecosystem, TDL lacks a principled and standardized framework for easily defining new architectures, restricting its accessibility and applicability. To address this issue, we introduce Generalized CCNNs (GCCNs), a simple yet powerful family of TDL models that can be used to systematically transform any (graph) neural network into its TDL counterpart. We prove that GCCNs generalize and subsume CCNNs, while extensive experiments on a diverse class of GCCNs show that these architectures consistently match or outperform CCNNs, often with less model complexity. In an effort to accelerate and democratize TDL, we introduce TopoTune, a lightweight software for defining, building, and training GCCNs with unprecedented flexibility and ease.
@@ 1 Introduction
To date, research in TDL has largely progressed by taking existing GNNs architectures (convolutional, attentional, message-passing, etc.) and generalizing them one-by-one to a specific TDL counterpart, whether that be on hypergraphs (Feng et al., 2019; Chen et al., 2020a; Yadati, 2020), on simplicial complexes (Roddenberry et al., 2021; Yang & Isufi, 2023; Ebli et al., 2020; Giusti et al., 2022a; Battiloro et al., 2024; Bodnar et al., 2021b; Maggs et al., 2024; Lecha et al., 2025), on cell complexes (Hajij et al., 2020; Giusti et al., 2022b; Bodnar et al., 2021a), or on combinatorial complexes (Battiloro et al., 2025; Eitan et al., 2024). Although overall valuable and insightful, such a fragmented research trend is slowing the development of standardized methodologies and software for TDL, as well as limiting the analysis of its cost-benefits trade-offs (Papamarkou et al., 2024). We argue these two challenges are hindering the use and application of TDL beyond the community of experts. This is particularly relevant as practitioners are beginning to turn to TDL for tackling application-specific scenarios, such as computer network modelling (Bernárdez et al., 2025).
@@
TopoX (Hajij et al., 2024b) and TopoBench(Telyatnikov et al., 2024) have become the reference Python libraries for developing and benchmarking TDL models, respectively. However, despite their potential in defining and implementing novel standardized methodologies in the field, the current focus of these packages is on replicating and analyzing existing message-passing CCNNs. Works like Jogl et al. (2022b, a) have instead focused on making TDL accessible by porting models to the graph domain. They do so via principled transformations from combinatorial topological domains to graphs. However, although these architectures are as expressive as their TDL counterparts (using the Weisfeiler-Lehman criterion (Xu et al., 2019b)), they are neither formally equivalent to nor a generalization of their TDL counterparts. Due to collapse of rank information during the graph expansion, the GNNs on the resulting graph do not preserve the same topological information.
@@
Contributions. This works seeks to accelerate TDL research and increase its accessibility and standardization for outside practitioners. To that end, we introduce a novel joint methodological and software framework that easily enables the development of new TDL architectures in a principled way—overcoming the limitations of existing works. We outline our main contributions and specify which of the field’s open problems (OPs), as defined in Papamarkou et al. (2024), they help answer:
@@
  • Systematic Generalization. We propose the first method to systematically generalize any neural network to its topological counterpart with minimal adaptation. Specifically, we define a novel expansion mechanism that transforms a combinatorial complex into a collection of graphs, enabling the training of TDL models as an ensemble of synchronized models. To our knowledge, this is the first method designed to accommodate many topological domains. (OPs 6, 11: foundational, cross-domain TDL.)
@@
  • General Architectures. Our method induces a novel wide class of TDL architectures, Generalized Combinatorial Complex Networks (GCCNs), portrayed in Fig. 1. GCCNs (i) formally generalize CCNNs, (ii) are cell permutation equivariant, and (iii) are as expressive as CCNNs. (OP 9: consolidating TDL advantages in a unified theory.)
@@
  • Implementation. We provide TopoTune, a lightweight PyTorch module for developing GCCNs fully integrated into TopoBench (Telyatnikov et al., 2024). (OP 4: need for software). Using TopoTune, practitioners can, for the first time, easily define and iterate upon TDL models, making TDL a much more practical tool for real-world datasets (OP 1: need for accessible TDL).
@@
  • Benchmarking. Using TopoTune, we create a broad class of GCCNs using four base GNNs and one base Transformer over two combinatorial topological spaces (simplicial and cell complexes). Unlike prior works that compare models under heterogeneous conditions, our systematic benchmarking provides a controlled evaluation of GCCNs across diverse architectures, datasets, and topological domains. A wide range of experiments on graph-level and node-level benchmark datasets shows GCCNs generally outperform existing CCNNs, often with smaller model sizes. Some of these results are obtained with GCCNs that cannot be reduced to standard CCNNs, further underlining our methodological contribution. A guide to the code is available at geometric-intelligence.github.io/topotune. (OP 3: need for standardized benchmarking.)
@@ 4 Generalized Combinatorial Complex Neural Networks
Proof of Prop. 4.1 (Appendix B.1) relies on setting the $\omega_{\mathcal{N}}$ of a GCCN to a simple, single-layer convolution.
@@
Proof of Prop 4.2 (Appendix B.2) hinges on the node-wise permutation equivariance of $\omega_{\mathcal{N}}$ and the permutation invariance of the inter-neighborhood aggregation.
@@
Proof of Prop. 4.3 (Appendix B.3) shows that GCCNs surpass CCNNs in expressivity by relating CCNNs to Weisfeiler-Leman (WL) and GCCNs to $k$-WL on augmented Hasse graphs.
Given Proposition 4.1, GCCNs allow us to define general TDL models using any neighborhood message function $\omega_{\mathcal{N}}$, such as any GNN. Not only does this framework avoid having to approximate CCNN computations, as is the case in previous works  [^4] These models employ GNNs running on one augmented Hasse graph, i.e. a GCCN that, given a collection of neighborhoods $\mathcal{N_{C}}$, uses a single neighborhood $\mathcal{N}_{tot}$ defined, for a cell $\sigma$, as $\mathcal{N}_{tot}(\sigma)=\bigcup_{\mathcal{N}\in\mathcal{N_{C}}}\mathcal{N}(\sigma)$. (Jogl et al., 2022b, a, 2023), but it also enjoys the same permutation equivariance as regular CCNNs (Proposition 4.2). We show in Appendix C that the resulting time complexity of a GCCN is a compromise between a typical GNN and a CCNN. Differently from the work in (Hajij et al., 2023), the fact that GCCNs can have arbitrary neighborhood message functions implies that non message-passing TDL models can be readily defined. For example, one could choose $\omega_{\mathcal{N}}$ to be a spectral convolution neural network such as Defferrard et al. (2016). To the best of our knowledge, GCCNs are the only objects in the literature that encompass all the above properties.
@@ 5 TopoTune
Our proposed methodology, together with its resulting GCCNs architectures, addresses the challenge of systematically generating principled, general TDL models. Here, we introduce TopoTune, a software module for defining and benchmarking GCCN architectures on the fly—a vehicle for accelerating TDL research. A quick start guide to the code and tutorial are provided at geometric-intelligence.github.io/topotune. This section details TopoTune’s main features.
@@
TopoTune is fully integrated into TopoBench (Telyatnikov et al., 2024), a comprehensive package offering a wide range of standardized methods and tools for TDL. Practitioners can access ready-to-use models, training pipelines, tasks, and evaluation metrics, including leading open-source models from TopoX (Hajij et al., 2024b). In addition, TopoBench features the largest collection of topological liftings currently available—transformations that map graph datasets into higher-order topological domains. Together, TopoBench and TopoTune organize the vast design space of TDL into an accessible framework, providing unparalleled versatility and standardization for practitionners.
@@ 6 Experiments
Table 1: Cross-domain, cross-task, cross-expansion, and cross-$\omega_{\mathcal{N}}$ comparison of GCCN architectures with top-performing CCNNs benchmarked on TopoBench (Telyatnikov et al., 2024). Best result is in bold and results within 1 standard deviation are highlighted blue. Experiments are run with 5 seeds. We report accuracy for classification tasks and MAE for regression.
@@
CCNN (Best Model on TopoBench) 80.43 ± 1.78 76.13 ± 2.70 76.67 ± 1.48 75.35 ± 1.50 0.34 ± 0.01 87.44 ± 1.28 75.63 ± 1.58 88.64 ± 0.36
@@
CCNN (Best Model on TopoBench) 76.17 ± 6.63 75.27 ± 2.14 76.60 ± 1.75 77.12 ± 1.07 0.36 ± 0.02 82.27 ± 1.34 71.24 ± 1.68 88.72 ± 0.50
@@
CCNN (Best Model on TopoBench) 80.43 ± 4.09 76.63 ± 1.74 75.18 ± 1.24 74.93 ± 2.50 0.51 ± 0.01 88.92 ± 0.44 74.93 ± 1.39 89.62 ± 0.25
We present experiments showcasing a broad class of GCCN’s constructed with TopoTune. These models consistently match, outperform, or finetune existing CCNNs, often with smaller model sizes. TopoTune’s integration into the TopoBench experiment infrastructure ensures a fair comparison with CCNNs from the literature, as data processing, domain lifting, and training are homogeonized.
@@ 6.1 Experimental Setup
We generate our class of GCCNs by considering ten possible choices of neighborhood structure $\mathcal{N}_{\mathcal{C}}$ (including both regular and per-rank, see Appendix E.1) and five possible choices of $\omega_{\mathcal{N}}$: GCN (Kipf & Welling, 2017), GAT (Velickovic et al., 2017), GIN (Xu et al., 2019a), GraphSAGE (Hamilton et al., 2017), and Transformer (Vaswani et al., 2017). We import these models directly from PyTorch Geometric (Fey & Lenssen, 2019) and PyTorch (Paszke et al., 2019). TopoTune enables running GCCNs on both an ensemble of strictly augmented Hasse graphs (eq. 6) and an augmented Hasse graph (eq. 5). While CCNN results reflect extensive hyperparameter tuning by Telyatnikov et al. (2024) (see that work’s Appendix C.2 for details), we largely fix GCCN training hyperparameters using the TopoBench default configuration.
@@
We include a wide range of benchmark tasks (see Appendix E.2). MUTAG, PROTEINS, NCI01, and NCI09 (Morris et al., 2020) are graph-level classification tasks about molecules or proteins. ZINC (Irwin et al., 2012) (subset) is a graph-level regression task about solubility. At the node level, the Cora, CiteSeer, and PubMed tasks (Yang et al., 2016) involve classifying publications within citation networks. We consider two topological domains: simplicial and cellular complexes. We use TopoBench’s lifting processes to infer higher-order relationships in these data. We use node features to construct edge and face features.
@@ 6.2 Results and Discussion
GCCNs outperform CCNNs.
Table 1 compares top-performing CCNNs with our class of GCCNs. GCCNs outperform CCNNs across all datasets in the simplicial and cellular domains and match hypergraph CCNNs—something CCNNs fail to achieve in node-level tasks. Across 16 domain/dataset combinations, GCCNs exceed the best CCNN by $>1\sigma$ in 11 cases. In 2 combinations, GCCNs are outperformed by the best GNN included in Telyatnikov et al. (2024). Representing complexes as ensembles of augmented Hasse graphs, rather than a single graph, improves results.
Generalizing existing CCNNs to GCCNs improves performance.
TopoTune makes it easy to iterate upon and improve preexisting CCNNs by replicating their architecture in a GCCN setting. For example, TopoTune can generate a counterpart GCCN by replicating a CCNN’s neighborhood structure, aggregation, and training scheme. We show in Table 2 that counterpart GCCNs can achieve comparable or better results than SCCN (Yang et al., 2022) and CWN (Bodnar et al., 2021a) just by sweeping over additional choices of $\omega_{\mathcal{N}}$. In the single augmented Hasse graph regime, GCCN models are consistently more lightweight, up to half their size (see Table 5).
Graph-Level Tasks Node-Level Tasks
Model MUTAG PROTEINS NCI1 NCI109 Cora Citeseer PubMed
SCCN (Yang et al., 2022)
CWN (Bodnar et al., 2021a)
GCCNs perform competitively to CCNNs with fewer parameters.
GCCNs are often more parameter efficient than existing CCNNs in simplicial and cellular domains, and in some instances (MUTAG, NCI1, NCI09), even in the hypergraph domain. We refer to Table 4. Even as GCCNs become more parameter-intensive for large graphs with high-dimensional embeddings—as seen in node-level tasks—they remain competitive. (We refer to Appendix H.1 for additional results on larger node-level datasets.) For instance, on the Citeseer dataset, a GCCN ($\omega_{\mathcal{N}}$ = GraphSAGE) outperforms the best existing CCNN while being 28% smaller. Training times in Appendix G show GCCNs train comparably on smaller datasets but slow down on larger ones, likely due to TopoTune’s on-the-fly graph expansion. Preprocessing this expansion in future work could mitigate the lag.
TopoTune finds parameter-efficient GCCNs.
By easily exploring a wide landscape of possible GCCNs for a given task, TopoTune helps identify models that maximize performance while minimizing model size. Fig. 5 illustrates this trade-off by comparing the performance and size of selected GCCNs (see Appendix I for more). On the PROTEINS dataset, two GCCNs using per-rank neighborhood structures (orange and black) achieve performance within 2% of the best result while requiring as little as 48% of the parameters. This reduction is due to fewer neighborhoods $\mathcal{N}$, resulting in fewer $\omega_{\mathcal{N}}$ blocks per GCCN layer. Similarly, on ZINC, lightweight neighborhood structures (orange and dark green) are competitive with reduced parameter costs. Node-level tasks, see less benefit, likely due to the larger graph sizes and higher-dimensional input features.
@@
Fig. 5 also provides insights into the impact of neighborhood message functions. On ZINC, GIN clearly outperforms all other models, which do not even appear in the plot’s range. In the less clear-cut cases of PROTEINS and Citeseer, we observe a trade-off between neighborhood structure and message function complexity. We find that larger base models (GIN, GraphSAGE) on lightweight neighborhood structures perform comparably to simpler base models (GAT, GCN) on larger neighborhood structures. This tradeoff warrants further research on the dataset-specific importance of neighborhood choice, or lack thereof. We refer to Appendix H.2 for additional experiments with more advanced GNNs, GATv2 (Brody et al., 2021) and PNA (Corso et al., 2020), as choices of $\omega_{\mathcal{N}}$.